how do you solve improper fractions ex.49 6/7+64 5/14
Answer by Theo(13342)

your equation is:

49 and 6/7 added to 64 and 5/14

this can be looked upon as:

49 + 6/7 + 64 + 5/14

set y = to this equation.

you get:

y = 49 + 6/7 + 64 + 5/14

multiply both sides of this equation by 14 to get

14y = 14*49 + 2*6 + 14*64 + 5

simplify and combine terms on the right side of this equation to get:

14y = 1599

divide both sides of this equation by 14 to get:

y = 1599/14 = 114.2142857

I confirmed this with the original equation so I know the answer is good.

The other way to solve this is to make the improper fractions into proper fractions and then solve.

your original equation is once again:

49 and 6/7 added to 64 and 5/14

take 49 and 6/7 and make a proper fraction out of it.

you multiply 49 by 7 and divide it by 7 to get 49*7/7

you add 49*7/7 to 6/7 to get (49*7+6)/7

this becomes 349/7

do the same with 64 and 5/14

multiply 64 by 14 and divide it by 14 to get 64*14/14.

this becomes 896/14.

add this to 5/14 to get 901/14

your problem becomes adding 349/7 to 901/14

you still need to get a common denominator so take 349/7 and multiply it by 2 and divide it by 2 to get 698/14.

your problem is now to add 698/14 to 901/14

this becomes (698 + 901)/14 which becomes 1599/14

divide 1599 by 14 to get 114.2142857

this is the same as 114 + 3/14

that's the same answer we got before so we're good.

the third alternative is to simply take the fractional part and turn it into a decimal and add it to the whole number part.

this works best if you have a calculator.

49 + 6/7 becomes 49.857142857

64 and 5/14 becomes 64.357142857

49.857142857 and 64.357142857 = 114.2142857 which is the same answer you got before using the two other methods.

when you multiply a number and divide a number by the same number, the number remains the same.

2 * 7/7 is still 2.

it just looks different.

it looks like 14/7 rather than 2.

if you divide 14 by 7 though, you still get 2.